### 29 Jan 2024 v1

# SelectLLM: Can LLMs Select Important Instructions to Annotate?

Abstract
Training large language models (LLMs) with a large and diverse instruction dataset aligns the models to comprehend and follow human instructions. Recent works have shown that using a small set of high-quality instructions can outperform using large yet more noisy ones. Because instructions are unlabeled and their responses are natural text, traditional active learning schemes with the model’s confidence cannot be directly applied to the selection of unlabeled instructions. In this work, we propose a novel method for instruction selection, called SelectLLM, that leverages LLMs for the selection of high-quality instructions. Our high-level idea is to use LLMs to estimate the usefulness and impactfulness of each instruction without the corresponding labels (i.e., responses), via prompting. SelectLLM involves two steps: dividing the unlabelled instructions using a clustering algorithm (e.g., CoreSet) to multiple clusters, and then prompting LLMs to choose high-quality instructions within each cluster. SelectLLM showed comparable or slightly better performance on the popular instruction benchmarks, compared to the recent state-of-the-art selection methods. All code and data are publicly available [^1] https://github.com/minnesotanlp/select-llm.
1 Introduction
Instruction-tuning, which fine-tunes language models (LLMs) to follow the instructions constructed from a massive number of diverse and different tasks, has shown impressive generalization performance on various unseen tasks (Wei et al., 2022; Chung et al., 2022). However, creating large and diverse annotated instruction datasets requires the huge cost of human labeling, and hence it becomes a major challenge for instruction-tuning. While synthetic datasets generated by advanced large language models (LLMs) like GPT-4 have partly addressed this issue (Taori et al., 2023; Wang et al., 2022a), they often contain low-quality data, highlighting the need for more focus on dataset refinement (Zhou et al., 2023; Cao et al., 2023).
Figure 1: Conceptual comparison between previous approaches to select instructions and SelectLLM. Focusing on input instructions (top) is unable to consider the difficulty or uncertainty of response. Response-based approaches (middle) can suffer from the inference cost and quality issues of synthetic responses. SelectLLM (bottom) does not suffer from these issues by estimating the effectiveness of instructions via prompting LLMs. More discussions are in Appendix B.
A typical solution to address these challenges is to allocate human resources, i.e., developing a way to select important instructions after being labeled, similar to active learning (Settles, 2009). Since it is relatively easy to access unlabelled instructions or texts from various sources, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. Previous selection algorithms usually consider density at the embedding space (Sener and Savarese, 2018) or uncertainty (or difficulty) of the generated predictions by the model (Kung et al., 2023); however, these methods are limited by being unable to incorporate difficulty of each sample or the cost of generating the inferences and quality issues from using synthetic labels, respectively (Figure 1). On the other hand, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we aim to explore a novel way to select unlabelled instructions by leveraging LLMs without inferences of their possible responses.
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. Our high-level idea is using LLMs to estimate the usefulness and impact of each instruction without the corresponding labels; through the initial experiments presented in Figure 3, we first verify that LLMs do possess such a capability. To enhance the effectiveness of instruction selection via LLMs, SelectLLM first divides the entire dataset into several small subsets to construct the input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving the global structure of the entire dataset. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these constructed queries into LLMs, and SelectLLM selects a few instructions that are expected to be the most helpful in fine-tuning LLMs.
We demonstrate the effectiveness of SelectLLM compared to various state-of-the-art selection methods on two popular benchmarks for instruction-tuning, Dolly (Conover et al., 2023) and Cleaned Alpaca (Taori et al., 2023). In the experiments, we verified that our method consistently outperforms the baselines across the varied selection sizes. For instance, when selecting samples from Dolly, SelectLLM exhibited nearly 2.5% - 3% relative improvements compared to the strongest baseline on the Rouge F1 score and cosine similarity, respectively across all sample sizes (1k/2k/3k). In addition, SelectLLM shows better cross-task generalizations, an important characteristic of instruction-tuned LMs. More interestingly, we observe that the generated responses by SelectLLM are preferred to the responses from other baselines when evaluated with GPT-4.
While the input prompt is designed to focus on improving the generic performance of LLMs, SelectLLM has flexibility that could be easily tuned for the specific purpose of the users, e.g., reducing the toxicity. Hence, we believe that our framework could open up the potential of LLMs for selecting the samples with any desired property, which could be flexibly set depending on the specific use case.
2 Related Work
Figure 2: Illustration of the proposed SelectLLM.
Instruction-tuning for LMs. Instruction-tuning (Wei et al., 2022), a form of fine-tuning LLMs, has emerged as a prominent methodology to align pre-trained LMs for various tasks by describing their descriptions into the common form of instruction. Due to its ease of implementation and remarkable generalization capabilities for unseen tasks (Wei et al., 2022; Chung et al., 2022; Jang et al., 2023), it has gained substantial popularity recently. Constructing these instruction datasets with human annotations is a standard way (Conover et al., 2023; Sanh et al., 2021; Wang et al., 2022b), but it has a certain difficulty in terms of variety of instructions and the total number of instances due to labeling cost (Wang et al., 2022a); One promising solution for this limitation is to synthesize existing datasets and create diverse, multi-task datasets with the help of LLMs such as Alpaca (Taori et al., 2023) and Self-instruct (Wang et al., 2022a). However, using LLM-creating data increases the risk of including low-quality examples, and it is known that removing such noise from the dataset is critical for the effective instruction-tuning of LLMs (Zhou et al., 2023; Cao et al., 2023); therefore, in this work, we explore an alternative way to use LLMs for constructing a high-quality instruction dataset, by using LLMs for selecting unlabeled instructions.
Sample selection for efficient instruction-tuning. Expanding instruction dataset comes with its own set of challenges, including resource and time constraints, the need for human annotation, and the prevalence of redundant data. Incorporating human efforts is one of the most straightforward ways to address these limitations; for example, Zhou et al. (2023) manually annotate and filter out the noisy ones using human annotators to select high-quality instructions. However, it could require huge costs when all the procedures are done by human labor; collecting abundant unlabelled instructions, and then selecting and annotating the useful ones could be an effective way to mitigate these issues, similar to the popular task of active learning (Settles, 2009; Sener and Savarese, 2018). Here, the most important thing is selection criteria.
One line of the works focuses on density to include diverse instructions; for example, Chen et al. (2023a) maps the unlabelled instructions into embedding space and conducts K-means clustering (Hartigan and Wong, 1979) and K-center greedy algorithm (Sener and Savarese, 2018) for the selection. Another line of the works focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) propose to measure the uncertainty by inducing the perturbation on the input instruction and see how the responses are changed. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to be useful in training LLMs; from LLMs’ capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. On the other hand, one recent work (Chen et al., 2023b) has used the prompting of LLMs to construct a high-quality instruction dataset. However, our work is totally different as we focus on selecting the unlabelled instructions while (Chen et al., 2023b) focus on filtering low-quality labeled instructions.
3 SelectLLM: Select Important Unlabelled Instructions using LLMs
3.1 Preliminary
We first describe the problem setup of our interest. Let $\mathcal{X}=\{x_{i}\}_{i=1}^{M}$ denote the given un-annotated dataset, where $x_{i}$ represents $i$th instruction example. Then, our goal is to select $N$ most effective instructions from $\mathcal{X}$ which will be labeled by human annotators, to fine-tune a given large language model (LLM) $f_{\theta}$, e.g., LLaMA-2 (Touvron et al., 2023), and make it be generalized for various instructions. Specifically, let’s assume that we select $N$ instructions under a criteria $s(j)$:
$\mathcal{X}(S)=\{x_{s(j)}\}_{j=1}^{N},$
$~{}\text{where}~{}S=\{s(j)|s(j)\in[1,M]\}_{j=1}^{N}$ (1)
where $S$ denotes the indices of selected instructions. Then, the selected instruction $x\in\mathcal{X}(S)$ is labeled with a corresponding response $y$ by human annotators, and it results in the annotated instruction dataset $\mathcal{D}$, where $\mathcal{D}=\{(x_{s(j)},y_{s(j)})\}_{j=1}^{N}$. 222 We assume that the human responses could be obtained right after the selection for the experiments. Therefore, the performance of fine-tuned LLMs on $\mathcal{X}(S)$ is significantly varied depending on which selection criteria $s(j)$ is considered. While various criteria have been explored under tasks like active learning (see Section 2), this direction is less explored under the regime of instruction-tuning.
Figure 3: Experiments to verify LLMs’ capability to infer the importance of unlabelled instructions. We prompt ChatGPT to sort the instructions based on their effectiveness for model training; then, we compare the performance of three fine-tuned LMs (LLaMA-2) on instructions with different ranks (First, center, and last). A full prompt is presented in Appendix C.
3.2 Selection via prompting LLMs
For the selection criteria $s(j)$, SelectLLM proposes to leverage LLMs with a properly designed prompt without using ground truth or generated responses. Our high-level intuition is that LLMs can infer the potential impact of each instruction by only reading the instruction; as shown in Figure 3, we observed that the recent LLMs, e.g., ChatGPT, could estimate the effectiveness of each instruction for model training (e.g., LLaMA-2), even without the corresponding responses. To further improve the effectiveness of selection via LLMs, we carefully designed the input prompt to incorporate several important perspectives for instruction-tuning, and it is presented in 4. Formally, this process could be described as follow: we first assume that the dataset $\mathcal{X}$ is divided into $K$ smaller and non-overlapped subsets, i.e., $\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{X}_{k}$. Then, we construct input query $q_{t}$ using the designed prompt $p_{\tt sel}$ and $\mathcal{X}_{k}$, and forward it to LLM to select $O=\lfloor N/K\rfloor$ examples:
$S_{t}=\texttt{LLM}\big{(}p_{\tt sel}(q_{t},O)\big{)}$ (2)
where $S_{t}=\{s_{t}(j)\}_{j=1}^{O},~{}s_{t}(j)\in[1,N]$.
The following are {N} candidate instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
Examine the provided list of {N} instructions, each uniquely identified by a number in brackets [].
Your task is to select {num} instructions that will be annotated by human annotators for model fine-tuning.
Look for instructions that are clear and relevant, exhibit a high level of complexity and detail, represent a diverse range of scenarios and contexts, offer significant instructional value and potential learning gain, and present unique challenges and specificity.
These selected instructions should ideally be the most beneficial for model fine-tuning after being annotated by human annotators.
Present your selections using the format []. e.g., [1,2] or [2,3].
The most impactful {num} instructions (only identifiers) are:
Figure 4: Designed input prompt of SelectLLM.
3.3 Composing query of LLMs via clustering
In addition, to further improve the effectiveness of using LLMs for selection, we carefully design how to divide the entire dataset into several subsets which would be used to construct input queries, based on the equal-size clustering method. Here, our high-level idea is composing the subsets that maximize the diversity among the instructions while maintaining the global structure of the dataset as much as possible. Specifically, we first extract the sentence embedding of the instructions in $\mathcal{X}$, using the pre-trained sentence encoder $g_{\phi}$, e.g., Sentence-BERT (Reimers and Gurevych, 2019). Then, we conduct K-means clustering (Hartigan and Wong, 1979) on these embeddings, and calculate $D\in\mathbb{R}^{N\times K}$, the distance of all instances in $\mathcal{X}$ to $K$ cluster centers $c_{1},\cdots,c_{K}$. Based on the distances, we assigned each instance $x$ among $[1,K]$, by iteratively taking the one with the shorted distance to the cluster center among the remaining instances, to guarantee equal sizes for each $k$.
Overall, the selection procedure of SelectLLM is as follows: (1) constructing the input queries by separating the entire dataset into multiple subsets of diverse instructions. Then, (2) feed these queries into LLM and get the selected indices. The formal presentation of these procedures is presented in Algorithms 1 and 2 in Appendix.
4 Experiments
Table 1: Experimental results on Dolly (Conover et al., 2023). Rouge (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.073 0.109 0.130 0.192 0.265 0.336 0.104 0.264 Perplexity 0.158 0.183 0.192 0.402 0.433 0.453 0.178 0.429 CBSsbert 0.147 0.200 0.216 0.359 0.473 0.512 0.188 0.448 Lengthlong 0.256 0.247 0.238 0.641 0.626 0.611 0.247 0.626 CBSinstr 0.258 0.255 0.255 0.617 0.638 0.632 0.256 0.629 Random 0.239 0.264 0.278 0.589 0.644 0.650 0.260 0.628 Diversity 0.237 0.275 0.282 0.582 0.650 0.666 0.265 0.633 OpenEnd 0.258 0.271 0.282 0.627 0.641 0.669 0.270 0.646 Coreset 0.271 0.281 0.279 0.649 0.662 0.659 0.277 0.657 Ours 0.278 0.288 0.289 0.668 0.680 0.686 0.285 0.678
In this section, we design our experiments to investigate the following questions:
  • $\circ$
    Does SelectLLM outperform the previous state-of-the-art selection methods for instruction-tuning LLMs? (Tables 1,  2, and 3)
  • $\circ$
    Do LLMs tuned with SelectLLM provide good responses qualitatively? (Tables 4 and 6)
  • $\circ$
    What is the effect of each component in SelectLLM? (Table 5)
  • $\circ$
    Why SelectLLM is more effective than other selection methods ? (Table 7)
4.1 Setups
Datasets and metrics. We use labeled datasets without using their responses/labels to test our hypothesis. Further, we utilize one human-generated dataset and one machine-generated dataset. For the former, we use (Conover et al., 2023) which is a combined effort of several Databricks employees, and as for the machine-generated dataset, we use Cleaned Alpaca which is based on (Taori et al., 2023) but with responses generated by GPT-4 and cleaned up to fix any errors in the prompts of the original dataset. When it comes to evaluating the performance, we analyze the similarity between the inferred text and the actual text using the Rouge L score. We also measure the cosine similarity between the inferred text and the actual text. Additionally, we perform a GPT-based analysis to gauge the effectiveness of the generated inferences.
Baselines. We consider several baselines for comparison with our algorithm as follows: (1) Random: selecting instances from the unlabeled dataset purely at random. (2) Length: Considers the length of input instruction, focusing on both longer and shorter ones to evaluate their impact (Lengthlong and Lengthshort). (3) Cluster-Based Selection (CBS) (Chen et al., 2023a): transforming instructions into embedding space, clustering them with HDBSCAN, and selecting samples using the K-Center-Greedy algorithm. We consider two different embedding spaces with Sentence-bert (Reimers and Gurevych, 2019) and Instructor (Su et al., 2023), and denote them as CBSsbert and CBSinstr, respectively. (4) Perplexity (Marion et al., 2023): selecting samples based on low per-token perplexity, indicating high model certainty and fluency. (5) Diversity (Wang et al., 2022a): for each instruction in the dataset, Rouge score is computed against a randomly selected subset comprising $n$ samples ($n\ll M$). Then, we select $k$ samples that exhibit the minimum Rouge scores. (6) Open-Endedness (OpenEnd) (Li et al., 2023): generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. (7) Coreset (Sener and Savarese, 2018): similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm (Sener and Savarese, 2018).
Table 2: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.219 0.263 0.261 0.519 0.632 0.625 0.248 0.592 Perplexity 0.264 0.278 0.272 0.628 0.646 0.640 0.271 0.638 CBSsbert 0.254 0.264 0.288 0.598 0.618 0.665 0.269 0.627 Lengthlong 0.228 0.266 0.297 0.550 0.618 0.683 0.264 0.617 CBSinstr 0.257 0.280 0.292 0.610 0.655 0.673 0.276 0.646 Random 0.281 0.281 0.271 0.653 0.662 0.656 0.278 0.657 Diversity 0.268 0.286 0.297 0.650 0.673 0.690 0.284 0.671 OpenEnd 0.250 0.247 0.276 0.616 0.601 0.645 0.258 0.621 Coreset 0.277 0.267 0.299 0.650 0.635 0.695 0.281 0.660 Ours 0.276 0.277 0.301 0.661 0.667 0.707 0.285 0.678
Implementation details. For our experiment, we utilize the Dolly and Cleaned Alpaca datasets for both the training and evaluation phases. Since there is no explicit train and test split for these datasets, we reserve 1k samples from each dataset to form our test sets. This allocation leaves us with 51k samples in the Cleaned Alpaca dataset and 14k samples in the Dolly dataset for training purposes. Then, we apply one of our sampling algorithms to each training dataset to select a subset of data, varying the subset size between 1k to 3k samples for training, with an 80:20 training and validation split. We fine-tune LLaMA-2 (7B) model (Touvron et al., 2023) by employing QLoRa (Dettmers et al., 2023), a model optimization technique, to reduce the memory requirements during the fine-tuning process. For the experiments, we use three different random seeds and then compute the average of the evaluation scores from these three models to derive a final score for each method. We run a total of 20 epochs with a batch size of 6. We use the Paged optimizer and set the gradient accumulation steps at 2. To avoid overfitting and select the best model, we integrate an Early Stopping Callback with a patience of 3 epochs and a threshold of 0.01. Also, for selecting instructions with SelectLLM, we commonly use ChatGPT (gpt-3.5-turbo-0613).
4.2 Main Results
In this section, we present our main experimental results. Our comprehensive evaluation, incorporating both Rouge (F1) scores and Cosine Similarity metrics, provides a detailed insight into the performance variations across different sample sizes (1k, 2k, and 3k), as well as the average performance across these sizes. The results for Dolly are detailed in Table 1 and for Alpaca in Table 2, respectively. To be specific, our analysis leads to several nuanced observations:
Dominant performance of SelectLLM. SelectLLM consistently outperforms other methods in the Dolly dataset, maintaining a lead with an average improvement of 2.6% in Rouge Score and 3% in Cosine Similarity across all sample sizes. This highlights SelectLLM’s adaptability and effectiveness in processing human-generated data. In the Cleaned Alpaca dataset, SelectLLM shows its strength particularly at the 1k and 3k sample sizes, outperforming others on the cosine similarity metric. While its performance at the 2k size is slightly lower, the overall trend underscores its reliability across various data volumes.
Consistent effectiveness across datasets. SelectLLM exhibits unparalleled consistency in both human and synthetic datasets. This uniformity across sample sizes sets it apart from other baselines and demonstrates its broad applicability. In contrast, other methods like Coreset, Diversity, and LengthBig show fluctuating performances depending on the dataset and sample size. For example, Coreset varies notably with sample size, while Diversity and LengthBig excel in the Cleaned Alpaca dataset but falter in the Dolly dataset. OpenEndedness, interestingly, performs better in Dolly but shows decreased effectiveness in Cleaned Alpaca. These variations in other methods further highlight the robust and adaptable nature of SelectLLM.
Table 3: Cross-dataset generalization. Rouge (F1) / Cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 3k examples on counterpart (e.g., train: Dolly $\rightarrow$ eval: Cleaned Alpaca) with different methods are compared. The best scores are highlighted in bold, and columns are the trained dataset.
Methods Dolly Cleaned Alpaca Random 0.205 / 0.589 0.260 / 0.669 OpenEnd 0.208 / 0.627 0.244 / 0.640 Coreset 0.208 / 0.651 0.271 / 0.684 SelectLLM 0.229 / 0.668 0.263 / 0.683
Cross-dataset generalization. We conducted an analysis to determine how well models trained on dolly samples using various sampling techniques generalize to cleaned alpaca data, and vice versa. Our findings are presented in Table 3 for both datasets. We made an intriguing observation with regards to models trained on dolly data - the model trained using our approach remarkably performed better than all the baselines by 10% on the cleaned alpaca test set. On the Cleaned Alpaca dataset, Coreset shows comparable performance to SelectLLM in terms of the Cosine Similarity and slightly better performance in rouge scores compared to SelectLLM. Further, the Cleaned Alpaca dataset appears to be a better dataset for cross-evaluation generalization when observing the performance of all baselines trained on it.
Evaluation with GPT-4. We evaluate the quality of generated responses between fine-tuned LLaMA-2 with SelectLLM and other methods with a 1k size dataset. Specifically, we randomly selected 1k size instructions, inputs from the original dataset, and generated responses from finetuned models. We ask GPT-4 to choose a better response to a given instruction, similar to (Liu et al., 2023). As shown in Table 4, when comparing SelectLLM to Random, SelectLLM won in 52% of the cases. We present the used prompt in Figure 6.
Table 4: Win-Tie-Draw from GPT-4 evaluation (%) on SelectLLM against Random and Coreset with Dolly.
Compared Methods Win Tie Lose SelectLLM vs Random 52.1 18.4 29.5 SelectLLM vs Coreset 44.2 25.4 30.4
Table 5: Ablation study. Rouge (F1) and cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 1k examples on Dolly with different methods are compared. The best scores are highlighted in bold.
[…]
Table 6: Examples of generated sentences from fine-tuned LLMs with different selection methods.
Instruction (#1): Where was Sarah Schechter born? Input: Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer. Ground Truth: Sarah Schechter was born in Boston, Massachusetts. Random Coreset Ours Sarah Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer.
### Instruction: normal-$\cdots$ ### Input: normal-$\cdots$ ### Response: normal-$\cdots$ ### Instruction:
Sarah Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer. Sarah Schechter was born in Boston, Massachusetts.
Table 7: Analysis of 1k selected instructions with different approaches on Dolly. Diversity is measured using kNN distance (Carbonera and Abel, 2015) with $k=1$. Perplexity is measured with GPT2-large (Radford et al., 2019), and Length is the number of characters.
Methods Diversity ($\uparrow$) Perplexity ($\downarrow$) Length Random 0.721 89 460 Coreset 0.931 47 847 OpenEnd 0.710 71 646 SelectLLM 0.796 30 1417
4.3 More Analyses with SelectLLM
4.3.1 Ablation study
To show the effectiveness of each component of SelectLLM, we ablate against different combinations of local selection prompting methods (Sec 3.2) and global division algorithms (Sec 3.3) on the 1k samples with Dolly.
Table 5 shows the results.
[…]
4.3.2 Qualitative results
Comparison of outputs from fine-tuned LLMs. Here, we compare the generated responses from fine-tuned LLMs using SelectLLM against two representative baselines, Random and Coreset. This comparison focuses on their adherence to and interpretation of instructions in both closed and open QA formats. The primary example of the analysis is detailed in Table 6, while additional examples are in the appendix (see Table 9).
In the example in Table 6, SelectLLM demonstrates superior precision and adherence to the instruction compared to the other methods. While the Coreset and Random methods provide correct responses, they tend to replicate the input text without much specificity. In contrast, SelectLLM generates a response that is not only precise and direct but also effectively extracts and succinctly presents the relevant information.
The additional examples, presented in Table 9, further underscore SelectLLM’s nuanced understanding and response capability. In an open-ended task, for instance, Random and Coreset deliver basic, concise responses, but lack the personalized and empathetic touches seen in SelectLLM’s reply. SelectLLM infuses its response with a sense of passion and commitment, demonstrating a deeper comprehension of the instruction’s intent. This sophistication in response is indicative of the advanced capabilities necessary in instruction-tuned models. Another closed QA example asking to differentiate between two terms shows all three models arriving at the correct answer, yet only SelectLLM clearly distinguishes between the two terms by providing detailed definitions, highlighting its capacity to deliver more informative and contextually rich responses.
Figure 5: Qualitative example of selection with a given query composed of 14 instructions on Dolly.
[…]
5 Conclusion
In this work, we present SelectLLM, a method to utilize LLMs for selecting an effective subset among unlabelled instructions. Through the various experiments based on two popular benchmarks, we validate the effectiveness of SelectLLM compared to the previous selection methods. This demonstrates its effectiveness in improving the efficiency of instruction-tuning for language models.
Limitations
Despite the impressive performance of SelectLLM, it is not without its limitations. A primary concern is the expense associated with utilizing LLMs like ChatGPT for data selection, which can be substantial. Additionally, the scalability of SelectLLM when dealing with exceptionally large datasets, or in scenarios requiring real-time data annotation, remains an area that needs further exploration. This aspect is particularly crucial given the ever-increasing size of datasets and the imperative for efficient processing in a wide range of practical applications. Hopefully, these limitations will be addressed in the future upon our work.
Broader Impact and Ethical Implications
The findings from our research not only establish the proficiency of LLMs in autonomously selecting high-quality data for training but also open new paths for future investigation in this domain. The successful application of LLMs in data-constrained environments is demonstrated by the exceptional ability of SelectLLM. This study, therefore, marks a significant stride in the field of Instruction-tuning for LLMs, paving the way for more efficient and effective training methodologies and expanding the scope of autonomous capabilities of LLMs.
Acknowledgements
The work has been supported by funding from Cisco. We would like to thank Cisco team and Minnesota NLP group members for providing us with valuable feedback and comments on the initial draft.
Appendix A More Details about Experiments
A.1 Datasets
1) Dolly (Conover et al., 2023): The Dolly dataset, developed by Databricks, Inc., is a comprehensive collection of over 15,000 human-generated instruction-following records. It encompasses a variety of behavioral categories such as brainstorming, classification, closed and open QA, generation, information extraction, and summarization. This dataset was created by Databricks employees and is designed to enhance the interactivity and responsiveness of large language models (LLMs) similar to ChatGPT.
2) Cleaned Alpaca: Cleaned Alpaca is based on (Taori et al., 2023) but with responses generated by GPT-4. It addresses several issues found in the original dataset, which was generated using GPT-3. The cleaned dataset incorporates approximately 52,000 instructions.
A.2 Baselines
1) Random: As the name suggests, random sampling involves selecting instances from the unlabeled dataset purely at random, without considering their informativeness or representativeness.
2) Coreset-Based Selection (Chen et al., 2023a): Uses clustering for sample selection. Involves transforming instructions into vector representations, clustering them with HDBSCAN, and selecting samples using the K-Center-Greedy algorithm which focuses on cluster centroids.
3) Perplexity (Li et al., 2023): Selects samples based on low per-token perplexity, indicating high model certainty and fluency, akin to the approach in (Li et al., 2023).
4) Diversity (Wang et al., 2022a): This method utilizes Rouge scores to evaluate the diversity of prompts. For each instruction in the dataset, Rouge scores are computed against a randomly selected subset comprising n samples, where n is strictly less than the total size of the dataset (n<Dataset Size). The selection criterion is based on the aggregation of these Rouge scores. Specifically, we select samples that exhibit the minimum aggregated Rouge scores, thereby ensuring a diverse representation in the final dataset.
5) Open-Endedness (Li et al., 2023): Determines prompt openness by generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. This process follows the open-endedness criteria defined for samples with a broader range of chain of thought reasonings in (Li et al., 2023).
[…]
7) Coreset (Sener and Savarese, 2018): Similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm.
8) Length: Considers the length of prompts, focusing on both longer and shorter ones to evaluate their impact.
A.3 SelectLLM
In Algorithms 1 and 2, we describe the proposed algorithms, presented in Section 3. In addition, we present examples of generated sentences from fine-tuned LLMs with different selection methods, in Table 9.
Question: Given the following responses to the target question, determine which is more informative and plausible to answer a given question properly.
Response 1:
{Method #1 response}
Response 2:
{Method #2 response}
Target Question:
{question}
Your Task:
Identify which response (Response 1 or Response 2) is more informative and plausible to answer a given question at hand. Choices: [Response 1, Response 2]. Answer with less than 3 words.
Answer:
Figure 6: Prompt for GPT-4 evaluation on SelectLLM against Random and Coreset. {blues} indicate the place for the inputs. To prevent order bias of LLMs, we ask GPT-4 twice with changed order of responses.
Algorithm 1 SelectLLM
Input: Un-annotated instructions $\mathcal{X}$, large language model LLM, input prompt $p_{\texttt{sel}}$, sentence encoder $g_{\phi}$, number of samples in query $K$, number of queries $T$, number of output $O$
Output: Selected indices $S_{\tt all}$
/* Construct input queries */
$q_{1},\cdots,q_{T}\leftarrow\texttt{Diverse-query}(\mathcal{X},g_{\phi},K,T)$
$S_{\tt all}\leftarrow\varnothing$
for $t=1$ to $T$ do
      /* Selection via LLM */
     $S_{t}\leftarrow\texttt{LLM}\big{(}p_{\tt sel}(q_{t},O)\big{)}$
     $S_{\tt all}\leftarrow S_{\tt all}\cup S_{t}$
end for
Algorithm 2 Diverse-query
Input: Un-annotated instructions $\mathcal{X}$, sentence encoder $g_{\phi}$, number of clusters $K$, number of queries $T$
Output: Set of queries $\{q_{t}\}_{t=1}^{T}$
$c_{1},\cdots,c_{K}\leftarrow\texttt{K-means}(\{g_{\phi}(x)\},K)$
$d_{1,1},\cdots,d_{N,K}\leftarrow\texttt{l2-dist}(\{g_{\phi}(x)\},\{c_{k}\})$
for $k=1$ to $K$ do
     $I_{k}\leftarrow\texttt{argsort}\{d_{1,k},\cdots,d_{N,k}\}$
end for
$A=\{1,\cdots,N\}$
for $t=1$ to $T$ do
     $q_{t}\leftarrow\varnothing$
     for $k=1$ to $K$ do
         $s\leftarrow 1$
         while $|q_{t}|<k$ do
              if $I_{k}(s)\in A$ then
                  $q_{t}\leftarrow q_{t}\cup\{x_{I_{k}(s)}\}$
              else
                  $s\leftarrow s+1$
              end if
         end while
     end for
end for
Appendix B Comparison with Previous Selection Methods
Table 8: Comparison of different selection methods.
Sources Complexity Information Flexibility Cost Input Low Low Low Low Response High High Low Medium LLMs Medium High High High
In this section, we provide a detailed comparison between SelectLLM and previous approaches for sample selection. First, we divide the existing approaches for sample selection into two different categories: input-based and response-based ones. Input-based approaches only use the input text to select samples, e.g., given instruction without the corresponding label. For example, Chen et al. (2023a) transforms input instruction into embedding space, and then applies clustering and k-center-greedy algorithms. In contrast, response-based approaches first generate responses with the external model, and then select samples using both instruction and artificial response; for instance, one can utilize the fine-tuned LLMs with small labeled instructions (Kung et al., 2023) or fixed pre-trained LLMs (Wang et al., 2023).
Since the two approaches rely on different sources to extract the information for the samples, they have distinct characteristics. First, in terms of the complexity of the method, an input-based one is much simpler than a response-based one as it does not require additional processes like model fine-tuning or generation of the response. However, response-based one can utilize more information about the sample, thanks to the generated response, while it requires more cost to obtain that. In the case of SelectLLM, it’s not very complex as the user can easily select the samples using the APIs. Also, SelectLLM utilizes extensive information within LLMs while inferring the importance of each sample to select the important ones with prompting. Although it requires the cost for prompting with APIs in our experiments, we remark that SelectLLM exhibits the unique capability that could be flexibly adapted for the desired property. We summarize the comparison of different approaches in Table 8.
Appendix C Verifying Capability of LLMs for Selecting Unlabelled Instruction
In this section, we present the full input prompt to verify whether LLMs could infer the importance of unlabelled instructions, which is presented in Figure 3. We adapted the prompt from the recent work using LLMs for text re-ranking (Sun et al., 2023) and the prompt is presented in Figure 7.
This is RankGPT, an intelligent assistant that can rank instructions based on their impactfulness and informativeness for model fine-tuning, when labeled by humans, like active learning.
The following are {num} examples of instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
I will rank the {num} instructions above based on their impactfulness and informativeness for model fine-tuning when labeled by humans, like active learning. The examples will be listed in descending order using identifiers, and the most impactful examples should be listed first, and the output format should be [] > [] > etc, e.g., [1] > [2] > etc.
The ranking results of the {num} examples (only identifiers) is
Figure 7: Full prompt to investigate the capability of LLM to infer the importance of instructions without labels.
Table 9: More examples of generated sentences.
Instruction (#2): Write an email to the head of academic coaching at Services for Students with Disabilities sharing that I’ve applied to the coaching position. I have experience volunteering with children on the autism spectrum. Ground Truth:
Hi,
I hope you are having a wonderful day. My name is _, and I just applied for the Program Staff position at the Services for Students with Disabilities to empower students who have disabilities to reach their academic dreams. My background working in higher education and volunteering with kids on the Autism spectrum combine to bring an empathetic and impactful perspective.
Please let me know if there is anything else I can do at this time to be considered for this position.
With gratitude,
_
Random Coreset Ours Dear Head of Academic Coaching,
I am writing to share that I have applied to the coaching position. I have experience volunteering with children on the autism spectrum.
Thank you for your time.
Sincerely,
Katie
I’m writing to share that I’ve applied to the coaching position. I have experience volunteering with children on the autism spectrum. Dear Head of Academic Coaching,
I am writing to share that I have applied to the coaching position. I have experience volunteering with children on the autism spectrum. I am passionate about helping children with disabilities and I believe that I would be a great asset to your team. I am confident that I can make a positive impact on the lives of the students that I work with.
Instruction (#3): Which is a species of fish? Red snapper or Red Sox Ground Truth: Red snapper Random Coreset Ours Red snapper
### Instruction:
What is the name of the first person to win the Tour de France?
### Input:
### Response:
Maurice Garin
### Instruction:
What is the name of the first person to win the Tour de France?
### Input:
### Response:
Maurice Garin
Red snapper is a species of fish. Red snapper is a species of fish.
Red Sox is a baseball team.
Appendix D More Analyses with Chosen Instructions
The following are {N} candidate instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
Examine the provided list of {N} instructions, each uniquely identified by a number in brackets [].
Your task is to select {num} instructions that will be annotated by human annotators for model fine-tuning.
Look for instructions that are clear and relevant, exhibit a high level of complexity and detail, represent a diverse range of scenarios and contexts, offer significant instructional value and potential learning gain, and present unique challenges and specificity.
These selected instructions should ideally be the most beneficial for model fine-tuning after being annotated by human annotators.
Present your selections using the format []. e.g., [1,2] or [2,3].
The most impactful {num} instructions (only identi-fiers) are: {prev_selection}
Explain why it was chosen, focusing on how it meets the above criteria and its potential contribution to model fine-tuning. Rationale for selection:
Figure 8: Prompt to generate reasoning for the selection.
In this section, we provide more details about additional analyses of chosen instructions. First, in Figure 8, we present a full prompt to generate reasoning for the selection by SelectLLM, which is used in Figure 5. We analyze 10 clusters, each containing 14 instructions, unveiling the LLM’s intricate selection criteria. Key factors influencing the LLM’s choices include clarity and relevance, complexity and detail, and the potential for instructional value and learning gains. The LLM shows a propensity for instructions that require a nuanced understanding and provide substantial learning opportunities, such as querying specific information about diverse topics like the Lollapalooza music festival, process mining, and the top speed of a Kia Stinger. Moreover, LLM consistently selects instructions that pose unique challenges and demand specificity, thereby testing and expanding its knowledge across various domains. Figures 9 and 10 show more such detailed examples with the rationales provided by the LLM for its selection compared against the selection by OpenEndedness.
Figure 9: Selection of an Instruction from a given cluster using LLM Based prompting (Red) along with its Chain of Thought Reasoning compared to selection based on the OpenEndedness scores of the given instructions (Blue).
Figure 10: Selection of an Instruction from a given cluster using LLM Based prompting (Red) along with its Chain of Thought Reasoning compared to selection based on the OpenEndedness scores of the given instructions (Blue). This is based on the example shown in the main paper.