### 05 Mar 2024 v4

# SelectLLM: Can LLMs Select Important Instructions to Annotate?

Abstract
Instruction tuning benefits from large and diverse datasets, however creating such datasets involves a high cost of human labeling. While synthetic datasets generated by large language models (LLMs) have partly solved this issue, they often contain low-quality data. One effective solution is selectively annotating unlabelled instructions, especially given the relative ease of acquiring unlabeled instructions or texts from various sources. However, how to select unlabelled instructions is not well-explored, especially in the context of LLMs. Further, traditional data selection methods, relying on input embedding space density, tend to underestimate instruction sample complexity, whereas those based on model prediction uncertainty often struggle with synthetic label quality. Therefore, we introduce SelectLLM, an alternative framework that leverages the capabilities of LLMs to more effectively select unlabeled instructions. SelectLLM consists of two key steps: Coreset-based clustering of unlabelled instructions for diversity and then prompting a LLM to identify the most beneficial instructions within each cluster. Our experiments demonstrate that SelectLLM matches or outperforms other state-of-the-art methods in instruction tuning benchmarks. It exhibits remarkable consistency across human and synthetic datasets, along with better cross-dataset generalization, as evidenced by a 10% performance improvement on the Cleaned Alpaca test set when trained on Dolly data. All code and data are publicly available [^1] https://github.com/minnesotanlp/select-llm.
1 Introduction
Instruction tuning, which fine-tunes language models (LMs) to follow instructions constructed from a massive number of diverse and different tasks, has shown impressive generalization performance on various unseen tasks (Wei et al., 2022; Chung et al., 2022). However, creating large and diverse annotated instruction datasets is a major challenge due to the significant cost of human labeling. While synthetic datasets generated by advanced large language models (LLMs) have partly addressed this issue (Taori et al., 2023; Wang et al., 2022a), they often contain low-quality data, highlighting the need for more focus on dataset refinement (Zhou et al., 2023; Cao et al., 2023).
Figure 1: Conceptual comparison between previous approaches to select instructions and SelectLLM. Focusing on input instructions (top) is unable to consider the difficulty or uncertainty of response. Output-based methods (middle) can suffer from the inference cost and quality issues of synthetic responses. SelectLLM (bottom) does not suffer from these issues by estimating the effectiveness of instructions via prompting LLMs.
As it is relatively easy to access unlabelled instructions or texts from various sources, one promising solution to address these challenges is developing a way to select important unlabelled instructions for annotation, similar to active learning (Settles, 2009); from this, one can create a high-quality small dataset if unlabelled instructions are well-selected and annotated accordingly. However, existing selection methods present limitations. Active learning-based algorithms focusing on density in the embedding space (Sener and Savarese, 2018) often overlook the complexity of instruction samples. Alternatively, output-based approaches that assess uncertainty in model predictions (Kung et al., 2023) grapple with high computational costs and the challenges posed by synthetic label quality (Figure 1). In light of these challenges, a few recent works show the capability of LLMs for measuring the quality or relevance of multiple texts (Sun et al., 2023; Liu et al., 2023). Inspired by these, we investigate whether LLMs can select effective unlabelled instructions, leveraging their vast knowledge base to discern the complexity and utility of each instruction without generating inferences.
Contribution. In this work, we introduce SelectLLM, a method that selects an effective subset of unlabelled instructions by prompting LLMs. At a high level, we use LLMs to estimate the usefulness and impact of each instruction without corresponding labels; through initial experiments presented in Figure 3, we first verify that LLMs possess such capability. To improve instruction selection via LLMs, SelectLLM first divides the dataset into several small subsets to construct input queries for prompting; specifically, we use equal-size K-means clustering to create each subset with diverse unlabelled instructions while preserving overall dataset structure. Then, SelectLLM constructs an input query for each subset using a carefully designed input prompt for selecting with LLMs. Finally, we forward these queries into the LLM, and SelectLLM selects a few instructions expected to be most helpful in fine-tuning models.
We demonstrate the effectiveness of SelectLLM compared to various state-of-the-art selection methods on two popular benchmarks for instruction tuning, Dolly (Conover et al., 2023) and Cleaned Alpaca (Taori et al., 2023). Our experiments show that SelectLLM consistently outperforms the baselines across the varied selection sizes. For instance, when selecting samples from Dolly, SelectLLM exhibited nearly 2.5% - 3% relative improvements compared to the strongest baseline on the Rouge-L F1 score and cosine similarity, respectively across all sample sizes (1k/2k/3k). In addition, SelectLLM shows better cross-task generalizations, an important characteristic of instruction-tuned LMs. More interestingly, we observe that the generated responses by SelectLLM are preferred to the responses from other baselines when evaluated with GPT-4.
Additionally, the input prompt in SelectLLM is primarily designed to select instructions that enhance the overall performance of LLMs. However, its flexible nature allows for easy customization to meet specific user needs (e.g., reducing toxicity), by simply adjusting SelectLLM’s input prompt during sample selection. This adaptability of our framework paves the way for using LLMs to select samples with various desired characteristics, tailored to suit distinct use cases, which is not possible with previous selection methods.
2 Related Work
Figure 2: Illustration of the proposed SelectLLM.
Instruction tuning for LMs. Instruction tuning (Wei et al., 2022), a form of fine-tuning LLMs, has emerged as a prominent methodology to align pre-trained LMs for various tasks by describing the tasks in the common form of instructions. Due to its ease of implementation and remarkable generalization capabilities for unseen tasks (Wei et al., 2022; Chung et al., 2022; Jang et al., 2023), it has gained substantial popularity recently. Constructing these instruction datasets with human annotations is a standard way (Conover et al., 2023; Sanh et al., 2021; Wang et al., 2022b), but this method faces challenges in terms of variety of instructions and the total number of instances due to labeling cost (Wang et al., 2022a); one promising solution for this limitation is to synthesize existing datasets and create diverse, multi-task datasets with the help of LLMs such as Alpaca (Taori et al., 2023) and Self-instruct (Wang et al., 2022a). However, using LLM-created data increases the risk of including low-quality examples, and it is known that removing such noise from the dataset is critical for the effective instruction tuning of LLMs (Zhou et al., 2023; Cao et al., 2023); therefore, in this work, we explore an alternative way to use LLMs to construct a high-quality instruction dataset, by using LLMs to select unlabeled instructions.
Sample selection for efficient instruction tuning. Expanding instruction datasets has its own set of challenges, including the need for extensive resources, time, human annotation, and the prevalence of redundant data. A common solution involves human intervention, such as the approach by Zhou et al. (2023), which manually annotates high-quality instructions after filtering out low-quality data. However, there are huge costs associated with such procedures done by human labor. An effective solution is to gather a large pool of unlabeled instructions and then selectively annotate the most useful ones, similar to the popular task of active learning (Settles, 2009; Sener and Savarese, 2018). Here, the most important thing is selection criteria.
One line of work focuses on density to include diverse instructions; for example, Chen et al. (2023a) map the unlabelled instructions into an embedding space and use K-means clustering (Hartigan and Wong, 1979) and the K-center greedy algorithm (Sener and Savarese, 2018) for the selection. The idea of clustering to select instruction data is also adopted in a vision-language domain (Wei et al., 2023). Another line of work focuses on uncertainty (or difficulty) of instructions measured with LLMs’ outputs; Kung et al. (2023) measure uncertainty by observing how LLM-generated responses vary with changes in the input instructions. Unlike these approaches, we directly prompt LLMs with unlabelled instructions, to select the few examples expected to help train the model; from LLM’s capability of reasoning and generating useful responses, we assume that they could infer the impactfulness of the unlabelled instructions. Meanwhile, a recent study by (Chen et al., 2023b) also prompts LLMs to construct a high-quality instruction dataset. However, our work differs as we focus on selecting unlabelled instructions while (Chen et al., 2023b) focuses on filtering out low-quality labeled ones.
3 SelectLLM: Select Important Unlabelled Instructions Using LLMs
3.1 Preliminary
We first describe the problem setup of our interest. Let $\mathcal{X}=\{x_{i}\}_{i=1}^{M}$ denote the given unlabeled dataset, where $x_{i}$ represents $i$th unlabeled instruction and $M$ is the total number of instructions. Then, our goal is to select $N$ most effective instructions from $\mathcal{X}$ which will be labeled by human annotators, to fine-tune a target large language model (LLM) $f_{\theta}$, e.g., LLaMA-2 (Touvron et al., 2023), and make it be generalized for various instructions. Formally, we select $N$ instructions from $\mathcal{X}$ under a selection criteria $s(j)$:
$\mathcal{X}(S)=\{x_{s(j)}\}_{j=1}^{N},$
$~{}\text{where}~{}S=\{s(j)|s(j)\in[1,M]\}_{j=1}^{N}$ (1)
where $S$ denotes the indices of selected instructions. Then, the selected instruction $x\in\mathcal{X}(S)$ is labeled with a corresponding label $y$ by human annotators, and it results in the annotated instruction dataset $\mathcal{D}$, where $\mathcal{D}=\{(x_{s(j)},y_{s(j)})\}_{j=1}^{N}$. Therefore, the performance of LLM $f_{\theta}$ fine-tuned on $\mathcal{X}(S)$ significantly varies depending on which selection criteria $s(j)$ is considered. While various selection criteria have been explored under tasks like active learning (see Section 2), this direction is less explored under the paradigm of instruction tuning.
Figure 3: Experiments to verify LLMs’ capability to infer the importance of unlabelled instructions. We prompt ChatGPT to sort the instructions based on their effectiveness for model training; then, we compare the performance of three fine-tuned LMs (LLaMA-2) on instructions with different ranks (First, center, and last). A full prompt is presented in Appendix C.
3.2 Selection via prompting LLMs
For the selection criteria $s(j)$, SelectLLM proposes to use LLMs with a properly designed prompt without using ground truths or generated labels. Our high-level intuition is that LLMs can infer the potential impact of each instruction by only reading the instruction; as shown in Figure 3, we observed that the recent LLM, e.g., ChatGPT, could estimate the effectiveness of each instruction for model training (e.g., LLaMA-2), even without the corresponding labels. To further improve the effectiveness of selection via LLMs, we carefully designed the input prompt to incorporate several important perspectives for instruction tuning, and it is presented in Figure 4. Formally, this process could be described as follows: we first assume that the dataset $\mathcal{X}$ is divided into $K$ non-overlapped subsets, i.e., $\mathcal{X}=\bigcup_{k=1}^{K}\mathcal{X}_{k}$. Then, we construct input query $q_{k}$ using the designed prompt $p_{\tt sel}$ and $\mathcal{X}_{k}$, and forward it to LLM to select $\widetilde{N}=\lfloor N/K\rfloor$ examples:
$S_{k}=\texttt{LLM}\big{(}p_{\tt sel}(q_{k},\widetilde{N})\big{)}$ (2)
where $S_{k}=\{s_{k}(j)\}_{j=1}^{\widetilde{N}},~{}s_{k}(j)\in[1,N]$.
The following are {N} candidate instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
Examine the provided list of {N} instructions, each uniquely identified by a number in brackets [].
Your task is to select {num} instructions that will be annotated by human annotators for model fine-tuning.
Look for instructions that are clear and relevant, exhibit a high level of complexity and detail, represent a diverse range of scenarios and contexts, offer significant instructional value and potential learning gain, and present unique challenges and specificity.
These selected instructions should ideally be the most beneficial for model fine-tuning after being annotated by human annotators.
Present your selections using the format []. e.g., [1,2] or [2,3].
The most impactful {num} instructions (only identifiers) are:
Figure 4: Designed input prompt of SelectLLM.
3.3 Composing query of LLMs via clustering
To further improve the effectiveness of using LLMs for selection, we carefully design how to divide the entire dataset into several subsets which would be used to construct input queries, based on the equal-size clustering method. Here, our high-level idea is composing the subsets that maximize the diversity among the instructions while maintaining the global structure of the dataset. Specifically, we first extract the embeddings of the instructions in $\mathcal{X}$, using the pre-trained sentence encoder $g_{\phi}$ such as Sentence-BERT (Reimers and Gurevych, 2019a). Then, we conduct K-means clustering (Hartigan and Wong, 1979) on these embeddings, and calculate $D\in\mathbb{R}^{N\times K}$, the distance of all instances in $\mathcal{X}$ to $K$ cluster centers $c_{1},\cdots,c_{K}$. Based on the distances, we assigned each instance $x$ among $[1,K]$, by iteratively taking the one with the shorted distance to the cluster center among the remaining instances, to guarantee equal sizes for each $k$.
Overall, the selection procedure of SelectLLM is as follows: (1) construct input queries by separating the entire dataset into multiple subsets of diverse instructions. Then, (2) feed these queries into a LLM and get the selected indices. The formal presentation of these procedures is presented in Algorithms 1 and 2 in Appendix.
4 Experiments
Table 1: Experimental results on Dolly (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge-L (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.073 0.109 0.130 0.192 0.265 0.336 0.104 0.264 Perplexity 0.158 0.183 0.192 0.402 0.433 0.453 0.178 0.429 CBSsbert 0.147 0.200 0.216 0.359 0.473 0.512 0.188 0.448 Lengthlong 0.256 0.247 0.238 0.641 0.626 0.611 0.247 0.626 CBSinstr 0.258 0.255 0.255 0.617 0.638 0.632 0.256 0.629 Random 0.239 0.264 0.278 0.589 0.644 0.650 0.260 0.628 Diversity 0.237 0.275 0.282 0.582 0.650 0.666 0.265 0.633 OpenEnd 0.258 0.271 0.282 0.627 0.641 0.669 0.270 0.646 Coreset 0.271 0.281 0.279 0.649 0.662 0.659 0.277 0.657 Ours 0.278 0.288 0.289 0.668 0.680 0.686 0.285 0.678
In this section, we design our experiments to investigate the following questions:
  • $\circ$
    Does SelectLLM outperform the previous state-of-the-art selection methods for instruction tuning LLMs? (Tables 123) If so, why? (Table 7)
  • $\circ$
    Do LLMs tuned with SelectLLM provide good responses qualitatively? (Tables 46)
  • $\circ$
    What is the effect of each component in SelectLLM? (Table 5)
4.1 Setups
Datasets and metrics. We use labeled datasets without using their responses to test our hypothesis. Further, we utilize one human-generated dataset and one machine-generated dataset. For the former, we use (Conover et al., 2023) which is a combined effort of several Databricks employees, and as for the latter, we use Cleaned Alpaca which is based on (Taori et al., 2023) but is cleaned up to fix any errors in the input prompts of the original dataset and has responses generated by GPT-4. For evaluating the performance, we assess the similarity between inferred and actual texts using Rouge scores and cosine similarity. Additionally, we perform a GPT-based analysis to gauge the effectiveness of the generated inferences.
Baselines. We consider several baselines for comparison with our algorithm as follows: (1) Random: selecting instances from the unlabeled dataset purely randomly. (2) Length: Considers the length of input instruction, focusing on both longer and shorter ones to evaluate their impact (Lengthlong and Lengthshort). (3) Cluster-Based Selection (CBS) (Chen et al., 2023a): transforming instructions into embedding space, clustering them with HDBSCAN (Campello et al., 2013), and selecting samples using the K-Center-Greedy algorithm. We consider two different embedding spaces with Sentence-BERT (Reimers and Gurevych, 2019a) and InstructOR (Su et al., 2023), and denote them as CBSsbert and CBSinstr, respectively. (4) Perplexity (Marion et al., 2023): selecting samples based on low per-token perplexity, indicating high model certainty and fluency. (5) Diversity (Wang et al., 2022a): for each instruction in the dataset, Rouge score is computed against a randomly selected subset comprising $n$ samples ($n\ll M$). Then, we select $k$ samples that exhibit the minimum Rouge scores. (6) Open-Endedness (OpenEnd) (Li et al., 2023): generating three inferences per prompt, counting unique bigrams, and selecting samples with the greatest variety of bigrams. (7) Coreset (Sener and Savarese, 2018): Similar to CBS, transforming instructions into embedding space with Sentence-bert, then selecting samples with K-Center-Greedy algorithm (Sener and Savarese, 2018).
Table 2: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 by different numbers of examples are compared. The best and second best scores are highlighted in bold and underline, respectively.
Rouge-L (F1) Cosine Similarity Avg Across Sizes Methods $1k$ $2k$ $3k$ $1k$ $2k$ $3k$ Avg Rouge Avg Cosine Lengthshort 0.219 0.263 0.261 0.519 0.632 0.625 0.248 0.592 Perplexity 0.264 0.278 0.272 0.628 0.646 0.640 0.271 0.638 CBSsbert 0.254 0.264 0.288 0.598 0.618 0.665 0.269 0.627 Lengthlong 0.228 0.266 0.297 0.550 0.618 0.683 0.264 0.617 CBSinstr 0.257 0.280 0.292 0.610 0.655 0.673 0.276 0.646 Random 0.281 0.281 0.271 0.653 0.662 0.656 0.278 0.657 Diversity 0.268 0.286 0.297 0.650 0.673 0.690 0.284 0.671 OpenEnd 0.250 0.247 0.276 0.616 0.601 0.645 0.258 0.621 Coreset 0.277 0.267 0.299 0.650 0.635 0.695 0.281 0.660 Ours 0.276 0.277 0.301 0.661 0.667 0.707 0.285 0.678
Implementation Details. We utilize the Dolly and Cleaned Alpaca datasets for training and evaluation. Due to the absence of predefined train-test splits in these datasets, we allocated 1k samples from each for testing, leaving 14k and 51k samples in the Dolly and Cleaned Alpaca datasets, respectively, for training. We employed sampling algorithms to select subsets of 1k to 3k samples from each train set. The LLaMA-2 (7B) model was fine-tuned using QLoRa (Touvron et al., 2023) to optimize memory usage during training. For instruction selection with SelectLLM, ChatGPT (gpt-3.5-turbo-0613) was utilized. To ensure robustness in our results, we conducted experiments using three different random seeds. The models’ evaluation scores were averaged to derive a final score for each method. Further details regarding the training process are provided in the Appendix.
4.2 Main Results
In this section, we present our main experimental results. Our comprehensive evaluation, incorporating both Rouge scores and Cosine Similarity metrics, provides a detailed insight into the performance variations across different sample sizes (1k, 2k, and 3k), as well as the average performance across these sizes. The results for Dolly are detailed in Table 1 and for Alpaca in Table 2, respectively. To be specific, our analysis leads to several nuanced observations:
Dominant performance of SelectLLM. SelectLLM consistently outperforms other methods in the Dolly dataset, maintaining a lead with an average improvement of 2.6% in Rouge Score and 3% in Cosine Similarity across all sample sizes. This highlights SelectLLM’s adaptability and effectiveness in processing human-generated data. In the Cleaned Alpaca dataset, SelectLLM shows its strength particularly at the 1k and 3k sample sizes, outperforming others on the cosine similarity metric. While its performance at the 2k size is slightly lower, the overall trend underscores its reliability across various data volumes.
Consistent effectiveness across datasets. SelectLLM exhibits unparalleled consistency in both human and synthetic datasets. This uniformity across sample sizes sets it apart from other baselines and demonstrates its broad applicability. In contrast, other methods like Coreset, Diversity, and Lengthlong show fluctuating performances depending on the dataset and sample size. For example, Coreset varies notably with sample size, while Diversity and Lengthlong excel in the Cleaned Alpaca dataset but falter in the Dolly dataset. OpenEndedness performs better in Dolly but shows decreased effectiveness in Cleaned Alpaca. This further highlights the robust and adaptable nature of SelectLLM.
Cross-dataset generalization. We analyze how well models trained on dolly samples using various sampling techniques generalize to cleaned alpaca data, and vice versa. Our results are presented in Table 3 for both datasets. We made an intriguing observation with regards to models trained on dolly data - the model trained using our approach remarkably performed better than all the baselines by 10% on the cleaned alpaca test set. On the Cleaned Alpaca dataset, Coreset shows comparable performance to SelectLLM in terms of Cosine Similarity and slightly better performance in Rouge scores. Further, Cleaned Alpaca appears to be a better dataset for cross-evaluation generalization when observing the performance of all baselines trained on it.
Table 3: Cross-dataset generalization for 3k sample size. Rouge-L (F1) / Cosine similarity of generated responses from LLaMA-2. The best scores are highlighted in bold, and column names indicate the dataset trained on.
Methods Dolly Cleaned Alpaca Random 0.205 / 0.589 0.260 / 0.669 OpenEnd 0.208 / 0.627 0.244 / 0.640 Coreset 0.208 / 0.651 0.271 / 0.684 SelectLLM 0.229 / 0.668 0.263 / 0.683
Evaluation with GPT-4. We evaluate the quality of generated responses between LLaMA-2 fine-tuned with SelectLLM and other methods on a 1k size instruction dataset, randomly sampled from the Dolly dataset. We ask GPT-4 to choose the better response to a given instruction, similar to (Liu et al., 2023). As shown in Table 4, SelectLLM wins in 52% of the cases when compared to Random sampling, and 44% cases when compared to Coreset sampling, further showcasing better inference quality of SelectLLM. The prompt for the evaluation is provided in the appendix (Figure 6).
Table 4: Win-Tie-Draw from GPT-4 evaluation (%) on SelectLLM against Random and Coreset with Dolly.
Compared Methods Win Tie Lose SelectLLM vs Random 52.1 18.4 29.5 SelectLLM vs Coreset 44.2 25.4 30.4
Table 5: Ablation study. Rouge and cosine similarity of generated responses from fine-tuned LLaMA-2 by selecting 1k examples on Dolly with different methods are compared. The best scores are highlighted in bold.
[…]
Table 6: Examples of generated sentences from fine-tuned LLMs with different selection methods.
Instruction (#1): Where was Sarah Schechter born? Input: Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer. Ground Truth: Sarah Schechter was born in Boston, Massachusetts. Random Coreset Ours Sarah Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer.
### Instruction: normal-$\cdots$ ### Input: normal-$\cdots$ ### Response: normal-$\cdots$ ### Instruction:
Sarah Schechter was born in Boston and grew up in Brookline, Massachusetts and New York City. Her father, Danny Schechter, a grandson of Russian-Jewish immigrant socialists, was a human rights activist, author, filmmaker and television producer. Sarah Schechter was born in Boston, Massachusetts.
Instruction (#2): Write an email to the head of academic coaching at Services for Students with Disabilities sharing that I’ve applied to the coaching position. I have experience volunteering with children on the autism spectrum. Ground Truth:
Hi,
I hope you are having a wonderful day. My name is _, and I just applied for the Program Staff position at the Services for Students with Disabilities to empower students who have disabilities to reach their academic dreams. My background working in higher education and volunteering with kids on the Autism spectrum combine to bring an empathetic and impactful perspective. Please let me know if there is anything else I can do at this time to be considered for this position. With gratitude,
Random Coreset Ours Dear Head of Academic Coaching,
I am writing to share that I have applied to the coaching position. I have experience volunteering with children on the autism spectrum.
Thank you for your time.
Sincerely,
Katie
I’m writing to share that I’ve applied to the coaching position. I have experience volunteering with children on the autism spectrum. Dear Head of Academic Coaching,
I am writing to share that I have applied to the coaching position. I have experience volunteering with children on the autism spectrum. I am passionate about helping children with disabilities and I believe that I would be a great asset to your team. I am confident that I can make a positive impact on the lives of the students that I work with.
Table 7: Analysis of 1k selected instructions with different approaches on Dolly. Diversity is measured using kNN distance (Carbonera and Abel, 2015) with $k=1$. Perplexity is measured with GPT2-large (Radford et al., 2019), and Length is the number of characters.
Methods Diversity ($\uparrow$) Perplexity ($\downarrow$) Length Random 0.721 89 460 Coreset 0.931 47 847 OpenEnd 0.710 71 646 SelectLLM 0.796 30 1417
4.3 More Analyses with SelectLLM
4.3.1 Ablation study
To show the effectiveness of each component of SelectLLM, we ablate against different combinations of local selection prompting methods (Sec 3.2) and global division algorithms (Sec 3.3) on 1k samples from Dolly. Table 5 shows the results.
[…]
Figure 5: Qualitative example of selection with a given query composed of 14 instructions on Dolly.
4.3.2 Qualitative results
Comparison of outputs from fine-tuned LLMs. We evaluate responses from LLMs fine-tuned with SelectLLM against Random and Coreset baselines, focusing on their handling of closed and open QA formats. Key examples are in Table 6, with more in the appendix (Table 9).
In the first example in Table 6, SelectLLM demonstrates superior precision and adherence to the instruction compared to the other methods. While the Coreset and Random methods provide correct responses, they tend to replicate the input text without much specificity. In contrast, SelectLLM generates a response that is not only precise and direct but also effectively extracts and succinctly presents the relevant information.
The second example, based on an Open-ended task further underscores SelectLLM’s nuanced understanding and response capability. While Random and Coreset give basic, concise answers, SelectLLM adds personalized, empathetic elements (highlighted in blue), showing deeper instruction comprehension. This reflects the advanced response capabilities needed in instruction-tuned models.
[…]
5 Conclusion
We introduce SelectLLM, a new approach that uses LLMs to choose an efficient subset from a set of unlabelled instructions. Our experiments on two popular benchmarks show that SelectLLM is more effective than previous selection methods. This demonstrates how LLMs can improve the efficiency of instruction tuning for language models.
Limitations
Despite the impressive performance of SelectLLM, it is not without its limitations. A primary concern is the expense associated with utilizing LLMs like ChatGPT for data selection, which can be substantial. Additionally, the scalability of SelectLLM when dealing with exceptionally large datasets, or in scenarios requiring real-time data annotation, remains an area that needs further exploration. This aspect is particularly crucial given the ever-increasing size of datasets and the imperative for efficient processing in a wide range of practical applications. Hopefully, these limitations will be addressed in the future upon our work.
Broader Impact and Ethical Implications
The findings from our research not only establish the proficiency of LLMs in autonomously selecting high-quality data for training but also open new paths for future investigation in this domain. The successful application of LLMs in data-constrained environments is demonstrated by the exceptional ability of SelectLLM. This study, therefore, marks a significant stride in the field of instruction tuning for LLMs, paving the way for more efficient and effective training methodologies and expanding the scope of autonomous capabilities of LLMs. In terms of ethical implications, the potential for any risk is limited to the application of LLMs in our framework, and the general risks associated with them such as LLMs showing bias in selecting certain instructions according to what it believes to be an impactful instruction. Further, bias can also be introduced based on how the prompt is designed by the user, when querying the LLMs in our framework.
Acknowledgements
This project has been generously supported by funding from Cisco. We thank the Cisco team and Minnesota NLP group members for their invaluable feedback on both the project and this draft.
Appendix A More Details about Experiments
A.1 Datasets
1) Dolly (Conover et al., 2023): The Dolly dataset, developed by Databricks, Inc., is a comprehensive collection of over 15,000 human-generated instruction-following records. It encompasses a variety of behavioral categories such as brainstorming, classification, closed and open QA, generation, information extraction, and summarization. This dataset was created by Databricks employees and is designed to enhance the interactivity and responsiveness of large language models (LLMs) similar to ChatGPT.
2) Cleaned Alpaca: Cleaned Alpaca is based on (Taori et al., 2023) but with responses generated by GPT-4. It addresses several issues found in the original dataset, which was generated using GPT-3. The cleaned dataset incorporates approximately 52,000 instructions.
A.2 Baselines
1) Random: As the name suggests, random sampling involves selecting instances from the unlabeled dataset purely at random, without considering their informativeness or representativeness.
[…]
3) Perplexity (Li et al., 2023): Selects samples based on low per-token perplexity, indicating high model certainty and fluency, akin to the approach in (Li et al., 2023).
4) Diversity (Wang et al., 2022a): This method utilizes Rouge scores to evaluate the diversity of prompts. For each instruction in the dataset, Rouge scores are computed against a randomly selected subset comprising n samples, where n is strictly less than the total size of the dataset (n < Dataset Size). The selection criterion is based on the aggregation of these Rouge scores. Specifically, we select samples that exhibit the minimum aggregated Rouge scores, thereby ensuring a diverse representation in the final dataset.
5) Open-Endedness (Li et al., 2023): Determines prompt open-endedness by generating three inferences per prompt, counting unique bigrams between the generated inferences, and selecting samples with the greatest variety of bigrams. This process follows the open-endedness criteria defined for samples with a broader range of chain of thought reasonings in (Li et al., 2023).
6) Coreset (Sener and Savarese, 2018): Similar to CBS, this method involves transforming data points into their embedding space with Sentence-BERT, then iterates a process to get subsets that maximize the coverage and diversity within each subset with a predetermined subset size. The algorithm selects samples by prioritizing those that maximize the distance to the nearest point already included in the subset, ensuring that the selected samples are diverse within each subset.
7) Length: Considers the length of prompts (Instruction + Input), focusing on both longer and shorter ones to evaluate their impact.
A.3 SelectLLM
In Algorithms 1 and 2, we describe the proposed algorithms, presented in Section 3. In addition, we present examples of generated sentences from fine-tuned LLMs with different selection methods, in Table 9.
A.4 Implementation details
For our experiment, we utilize the Dolly and Cleaned Alpaca datasets for both the training and evaluation phases. As there were no explicit train and test split for these datasets, we randomly sampled 1k samples from each dataset to form our test sets. This allocation leaves us with 51k samples in the Cleaned Alpaca dataset and 14k samples in the Dolly dataset for training purposes. Then, we apply one of our sampling algorithms to each training dataset to select a subset of data, varying the subset size between 1k to 3k samples for training, with an 80:20 training and validation split. We fine-tune LLaMA-2 (7B) model (Touvron et al., 2023) by employing QLoRa (Dettmers et al., 2023), a model optimization technique, to reduce the memory requirements during the fine-tuning and inference processes. For the experiments, we use three different random seeds and then compute the average of the evaluation scores from these three models to derive a final score for each method. We run a total of 20 epochs with a batch size of 6. We use the Paged optimizer and set the gradient accumulation steps at 2. To avoid overfitting and select the best model, we integrate an Early Stopping Callback with a patience of 3 epochs and a threshold of 0.01. Also, for selecting instructions with SelectLLM, we commonly use ChatGPT (gpt-3.5-turbo-0613).
Question: Given the following responses to the target question, determine which is more informative and plausible to answer a given question properly.
Response 1:
{Method #1 response}
Response 2:
{Method #2 response}
Target Question:
{question}
Your Task:
Identify which response (Response 1 or Response 2) is more informative and plausible to answer a given question at hand. Choices: [Response 1, Response 2]. Answer with less than 3 words.
Answer:
Figure 6: Prompt for GPT-4 evaluation on SelectLLM against Random and Coreset. {blues} indicate the place for the inputs. To prevent order bias of LLMs, we ask GPT-4 twice with changed order of responses.
Algorithm 1 SelectLLM
Input: Un-annotated instructions $\mathcal{X}$, large language model LLM, input prompt $p_{\texttt{sel}}$, sentence encoder $g_{\phi}$, number of samples in query $K$, number of queries $T$, number of output $O$
Output: Selected indices $S_{\tt all}$
/* Construct input queries */
$q_{1},\cdots,q_{T}\leftarrow\texttt{Diverse-query}(\mathcal{X},g_{\phi},K,T)$
$S_{\tt all}\leftarrow\varnothing$
for $t=1$ to $T$ do
      /* Selection via LLM */
     $S_{t}\leftarrow\texttt{LLM}\big{(}p_{\tt sel}(q_{t},O)\big{)}$
     $S_{\tt all}\leftarrow S_{\tt all}\cup S_{t}$
end for
Algorithm 2 Diverse-query
Input: Un-annotated instructions $\mathcal{X}$, sentence encoder $g_{\phi}$, number of clusters $K$, number of queries $T$
Output: Set of queries $\{q_{t}\}_{t=1}^{T}$
$c_{1},\cdots,c_{K}\leftarrow\texttt{K-means}(\{g_{\phi}(x)\},K)$
$d_{1,1},\cdots,d_{N,K}\leftarrow\texttt{l2-dist}(\{g_{\phi}(x)\},\{c_{k}\})$
for $k=1$ to $K$ do
     $I_{k}\leftarrow\texttt{argsort}\{d_{1,k},\cdots,d_{N,k}\}$
end for
$A=\{1,\cdots,N\}$
for $t=1$ to $T$ do
     $q_{t}\leftarrow\varnothing$
     for $k=1$ to $K$ do
         $s\leftarrow 1$
         while $|q_{t}|<k$ do
              if $I_{k}(s)\in A$ then
                  $q_{t}\leftarrow q_{t}\cup\{x_{I_{k}(s)}\}$
              else
                  $s\leftarrow s+1$
              end if
         end while
     end for
end for
Appendix B Comparison with Previous Selection Methods
Table 8: Comparison of different selection methods.
Sources Complexity Information Flexibility Cost Input Low Low Low Low Response High High Low Medium LLMs Medium High High High
In this section, we provide a detailed comparison between SelectLLM and previous approaches for sample selection. First, we divide the existing approaches for sample selection into two different categories: input-based and response-based ones. Input-based approaches only use the input text to select samples, e.g., given instruction without the corresponding label. For example, Chen et al. (2023a) transforms input instruction into embedding space, and then applies clustering and K-Center-Greedy algorithms. In contrast, response-based approaches first generate responses with the external model, and then select samples using both instruction and artificial response; for instance, one can utilize the fine-tuned LLMs with small labeled instructions (Kung et al., 2023) or fixed pre-trained LLMs (Wang et al., 2023).
Since the two approaches rely on different sources to extract the information for the samples, they have distinct characteristics. First, in terms of the complexity of the method, an input-based one is much simpler than a response-based one as it does not require additional processes like model fine-tuning or generation of the response. However, response-based one can utilize more information about the sample, thanks to the generated response, while it requires more cost to obtain that. In the case of SelectLLM, it’s not very complex as the user can easily select the samples using the APIs. Also, SelectLLM utilizes extensive information within LLMs while inferring the importance of each sample to select the important ones with prompting. Although it requires the cost for prompting with APIs in our experiments, we remark that SelectLLM exhibits the unique capability that could be flexibly adapted for the desired property. We summarize the comparison of different approaches in Table 8.
Appendix C Verifying Capability of LLMs for Selecting Unlabelled Instruction
In this section, we present the full input prompt to verify whether LLMs could infer the importance of unlabelled instructions, which is presented in Figure 3. We adapted the prompt from the recent work using LLMs for text re-ranking (Sun et al., 2023) and the prompt is presented in Figure 7.
This is RankGPT, an intelligent assistant that can rank instructions based on their impactfulness and informativeness for model fine-tuning, when labeled by humans, like active learning.
The following are {num} examples of instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
I will rank the {num} instructions above based on their impactfulness and informativeness for model fine-tuning when labeled by humans, like active learning. The examples will be listed in descending order using identifiers, and the most impactful examples should be listed first, and the output format should be [] > [] > etc, e.g., [1] > [2] > etc.
The ranking results of the {num} examples (only identifiers) is
Figure 7: Full prompt to investigate the capability of LLM to infer the importance of instructions without labels.
Table 9: Another instance of a closed QA example asking to differentiate between two terms. All three models arrive at the correct answer, however only SELECTLLM clearly distinguishes between the two terms by providing detailed definitions, highlighting its capacity to deliver more informative and contextually rich explanations
Instruction (#3): Which is a species of fish? Red snapper or Red Sox Ground Truth: Red snapper Random Coreset Ours Red snapper
### Instruction:
What is the name of the first person to win the Tour de France?
### Input:
### Response:
Maurice Garin
### Instruction:
What is the name of the first person to win the Tour de France?
### Input:
### Response:
Maurice Garin
Red snapper is a species of fish. Red snapper is a species of fish.
Red Sox is a baseball team.
Table 10: Experimental results on Dolly (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 models on the entire dolly dataset are compared with a model trained on only 3k samples chosen by SelectLLM.
Methods Rouge-L (F1) Cosine Similarity Dolly-14k 0.292 0.689 SelectLLM-3k 0.289 0.686
Table 11: Experimental results on Cleaned Alpaca (Conover et al., 2023). Rouge-L (F1) and Cosine similarity of generated responses from fine-tuned LLaMA-2 models on the entire Cleaned Alpaca dataset is compared with a model trained on only 3k samples chosen by SelectLLM.
Methods Rouge-L (F1) Cosine Similarity CleanedAlpaca-51k 0.356 0.755 SelectLLM-3k 0.301 0.707
Appendix D More Analyses with Chosen Instructions
The following are {N} candidate instructions that describe a task, each indicated by a number identifier [].
[1]
### Instruction: {Example #1 Instruction}
### Input: {Example #1 Input}
[N]
### Instruction: {Example #N Instruction}
### Input: {Example #N Input}
Examine the provided list of {N} instructions, each uniquely identified by a number in brackets [].
Your task is to select {num} instructions that will be annotated by human annotators for model fine-tuning.
Look for instructions that are clear and relevant, exhibit a high level of complexity and detail, represent a diverse range of scenarios and contexts, offer significant instructional value and potential learning gain, and present unique challenges and specificity.
These selected instructions should ideally be the most beneficial for model fine-tuning after being annotated by human annotators.
Present your selections using the format []. e.g., [1,2] or [2,3].
The most impactful {num} instructions (only identi-fiers) are: {prev_selection}
Explain why it was chosen, focusing on how it meets the above criteria and its potential contribution to model fine-tuning. Rationale for selection:
Figure 8: Prompt to generate reasoning for the selection.
In this section, we provide more details about additional analyses of chosen instructions. First, in Figure 8, we present a full prompt to generate reasoning for the selection by SelectLLM, which is used in Figure 5. We analyze 10 clusters, each containing 14 instructions, unveiling the LLM’s intricate selection criteria. Key factors influencing the LLM’s choices include clarity and relevance, complexity and detail, and the potential for instructional value and learning gains. The LLM shows a propensity for instructions that require a nuanced understanding and provide substantial learning opportunities, such as querying specific information about diverse topics like the Lollapalooza music festival, process mining, and the top speed of a Kia Stinger. Moreover, LLM consistently selects instructions that pose unique challenges and demand specificity, thereby testing and expanding its knowledge across various domains. Figures 9 and 10 show more such detailed examples with the rationales provided by the LLM for its selection compared against the selection by OpenEndedness.
Appendix E Comparison of Selected Sample with the Entire Dataset
We compared the SelectLLM-3k sampled model with models trained on full Dolly and Cleaned Alpaca datasets. As detailed in 10 and 11, SelectLLM-3k nearly equals the full Dolly dataset model in performance, and achieves 85% of the full dataset’s performance on the Cleaned Alpaca dataset, as per the Rouge score. This underscores SelectLLM’s efficiency in filtering out noise from the entire dataset, especially in the case of Dolly. While our results are akin to the data filtering approach in (Chen et al., 2023b), which also samples 3k instructions from the Dolly dataset, there are notable differences. Their method outperforms the full dataset, possibly due to their inclusion of sample outputs in the filtering process, an aspect not considered in our approach.
Figure 9: Selection of an Instruction from a given cluster using LLM Based prompting (Red) along with its Chain of Thought Reasoning compared to selection based on the OpenEndedness scores of the given instructions (Blue).
Figure 10: Selection of an Instruction from a given cluster using LLM Based prompting (Red) along with its Chain of Thought Reasoning compared to selection based on the OpenEndedness scores of the given instructions (Blue). This is based on the example shown in the main paper.